# Sonja Hohloch  &\&&  Pedro Santos

### v1

@@ Abstract.
We adapt and generalize McDuff’s method of probes from toric system to so called semitoric integrable systems and apply it to study the (non)displaceability properties of the fibers of $3$ examples of semitoric integrable systems.
@@ 1. Introduction
Let $(M,\omega)$ be a closed symplectic manifold. The Arnold conjecture, formulated in the $1960$s, can be stated in many different ways: for instance via the fixed points of a Hamiltonian diffeomorphism $\phi$ or equivalently looking at the intersection points of the graph of $\phi$ and the diagonal in $M\times M$. When one endows $M\times M$ with the symplectic form $\omega\oplus-\omega$ the graph of $\phi$ and the diagonal become Lagrangian submanifolds. Therefore the question about the existence of fixed points is turned into a so called Lagrangian intersection problem. This was the approach taken by Floer [floer1988morse, floer1989witten, floer1988unregularized] towards the end of the 1980s to prove the Arnold conjecture under certain assumptions. Roughly speaking, for a Lagrangian submanifold $L\subset M$ with $\pi_{2}(M,L)=\{0\}$, Floer developed a homology theory generated by the intersection points of $L\cap\phi(L)$ where $\phi$ is a Hamiltonian diffeomorphism.
@@
Therefore, the interest in (non)displaceability is at least twofold: on the one hand motivated by the computation of Lagrangian Floer homology and on the other hand by symplectic rigidity. Over the years, research on (non)displaceability developed into an area of its own, for instance, just to mention a few, it is essential for Fukaya categories (see Fukaya & Oh & Ohta & Ono [fukayabook1, fukayabook2]), it motivated the development of symplectic quasi-states by Entov & Polterovich [polterovich2006quasi], and the development of explicit methods, see for example McDuff [mcduff2011displacing] and Abreu & Borman & McDuff [abreu2014displacing], that give sufficient conditions for a Lagrangian torus fiber to be displaceable. This last aspect is in fact the starting point for the present paper, as we will explain now.
Briefly summarized, McDuff’s [mcduff2011displacing] explicit method (called the method of probes) makes use of the Delzant polytope (= momentum polytope) of a symplectic toric manifold. In the world of integrable systems, toric systems are the simplest examples and are classified by their corresponding Delzant polytope. The next natural, more general class of systems are so called semitoric systems (see Definition 2.17) which were classified by Pelayo & Vũ Ngǫc [pelayo2009semitoric, pelayo2011constructing] and Palmer & Pelayo & Tang [palmer2019semitoric] in terms of $5$ invariants. One of these invariants, the so called polytope invariant, is the generalization of the Delzant polytope of a toric system. In the present paper we will generalize the method of probes from toric system to semitoric systems and apply it to several examples.
@@ 1.1. Our results
Let $(M,\omega,F=(J,H))$ be a semitoric system (see Definition 2.17) on a $4$-dimensional symplectic manifold. Intuitively semitoric systems are more general than toric systems by requiring only for an $S^{1}$-symmetry instead of a $\mathbb{T}^{2}$-symmetry and allowing for focus-focus singularities in addition to the ones appearing in the toric case. As in the toric situation the fibers are connected. A fiber with $k$ focus-focus points is often seen as a torus with $k$ pinches. Alternatively the fiber $F^{-1}(c)$ can be seen as a closed chain of Lagrangian $k$ spheres joint at the poles. If $k=1$ the sphere is immersed, and if $k>1$ the spheres are embedded. Our first result is about the (non)displaceability of focus-focus fibers with more than one focus-focus point:
@@
Figure 1.2. $A(f_{\vec{\epsilon}}(F(M)))$ is the result of applying a suitable integral affine transformation $A$ to the representative of the polytope invariant $f_{\vec{\epsilon}}(F(M))$. Let $0<h<a<\frac{R}{2}$. The square $[0,a]\times[0,a]$ fits inside the rectangle $[0,R]\times[0,a]$. The rectangle is inside $A(f_{\vec{\epsilon}}(F(M)))$ and does not intersect the cut associated with the focus-focus value $c_{2}$. The rectangle $[0,R]\times[0,a]$ represents the symplectic embedding of the product of disks $D^{2}(R)\times D^{2}(a)$. Under these conditions the focus-focus fiber given by $c_{1}$ is displaceable, since we can displace $D^{2}(a)\times D^{2}(a)$ inside $D^{2}(R)\times D^{2}(a)$ due to the fact that $a<\frac{R}{2}$ and we can "torically smoothen" the corner introduced by $c_{1}$.
@@
First, in Section 4.1 we study the coupled spin oscillator which is a semitoric system $(M,\omega,F)$ on $M=S^{2}\times\mathbb{R}^{2}$, for which the main result is:
@@
In Section 4.2 we study the coupled angular momenta and its variations. Let $R_{1},R_{2}\in\mathbb{R}^{>0}$ and $t\in[0,1]$ a parameter. Consider the product manifold $M:=S^{2}\times S^{2}$ with symplectic form $\omega:=-(R_{1}\omega_{S^{2}}\bigoplus R_{2}\omega_{S^{2}})$ where $\omega_{S^{2}}$ is the standard symplectic form on $S^{2}$. The coupled angular momenta system is a family of $4$-dimensional completely integrable systems $(M,\omega,F_{t}:=(L,H_{t}))$, with $R_{1}<R_{2}$. The case $R_{1}>R_{2}$ is called the reverse coupled angular momenta and the case $R_{1}=R_{2}$ is called the Kepler problem. For an appropriate range of the parameter $t$ a focus-focus point always exists.
@@
Note that the symplectic rigidity problem is sensitive to the choice of $R_{1}$ and $R_{2}$. For the Kepler problem concerning the question of (non)displaceability we get a different behavior:
@@
Let $R_{1}=R_{2}$ and $t$ be in the parameter range such that a focus-focus point exists. There exists a $t_{0}$ such that:
  • If $t<t_{0}$ there exists only a single nondisplaceable fiber. In particular the focus-focus fiber is displaceable.
@@
In Section 4.3 we start studying the (non)displaceability of the fibers of the toric system induced by the octagon $\Delta$ with vertices
@@
as a Delzant polytope. Then we consider the semitoric perturbation of this system, introduced in De Meulenaere & Hohloch [de2021family], and study the (non)displaceability properties of the semitoric fibers of this perturbation. Let $(M,\omega,F_{t})$ be the semitoric perturbation of the toric system. For $t=\frac{1}{2}$ the system has two focus-focus fibers with two focus-focus points and otherwise $4$ focus-focus fibers with a single focus-focus point. In Appendix A we show that a representative of the polytope invariant, for $t\neq\frac{1}{2}$, is given by the octagon $\Delta$.
@@
      The focus-focus fibers and the fiber over the point $(\frac{3}{2},\frac{3}{2})$ in any representative of the polytope invariant are nondisplaceable.
@@
      If the image of the focus-focus values in the representative of the polytope invariant given by the octagon are contained in the region $\{y>2\}\cap\{y<1\}$ then the fibers over the points $\{(1,1),(2,1),(\frac{3}{2},\frac{3}{2}),(1,2),(2,2)\}$ are nondisplaceable. All other fibers are displaceable. In particular the focus-focus fibers are displaceable.
@@ 1.2. Structure of the paper
  • In Section 4 we study the (non)displaceability properties of $3$ semitoric systems, namely the coupled spin oscillator, the coupled angular momenta and the semitoric system induced by the octagon.
@@ 2.2. Integrable systems
[…]
@@
Two integrable systems $(M,\omega,F)$ and $(M^{\prime},\omega^{\prime},F^{\prime})$ are isomorphic if there exists a diffeomorphism $\psi:M\rightarrow M^{\prime}$ such that $\psi^{*}\omega^{\prime}=\omega$ and a diffeomorphism $g:F(M)\rightarrow F^{\prime}(M^{\prime})$ such that $F^{\prime}\circ f=g\circ F$.
@@
A point $m\in M$ in which $X_{f_{1}}(m),...,X_{f_{n}}(m)$ of $T_{m}M$ are linearly dependent is said to be singular or a singularity of $F$. Otherwise we call it regular. The rank of $m$ is given by $\dim\operatorname{span}_{\mathbb{R}}\{X_{f_{1}}(m),...,X_{f_{n}}(m)\}$. If a fiber $F^{-1}(c)$ for some $c\in{\mathbb{R}}^{m}$ contains a singularity we call it singular and otherwise regular. Denote a $k$-dimensional torus briefly by $\mathbb{T}^{k}$. By the Liouville-Arnold-Mineur Theorem, see Arnold [arnol2013mathematical], a pre-image of a regular value is diffeomorphic to $\mathbb{T}^{n}$. Moreover, $\omega$ vanishes along regular fibers $F^{-1}(c)$, hence $F^{-1}(c)$ is a Lagrangian submanifold, traditionally called a Liouville torus. One often refers to an integrable system $(M,\omega,F)$ as a singular Lagrangian fibration where singular emphasizes that $F$ may have singular fibers of various kinds: for example tori of dimension $m\in\{0,...,n-1\}$, pinched-tori etc.
@@ 2.3. Action-angle coordinates
One important theorem about the symplectic geometry of integrable systems is the existence of so-called action-angle coordinates, see Arnold [arnol1963theorem]: each regular fiber, in addition to being diffeomorphic to $\mathbb{T}^{n}$, can be seen as sitting inside of the cotangent bundle $T^{*}\mathbb{T}^{n}$ as the zero section and, in a neighborhood of the fiber, the integrable system can be written in a symplectic normal form $F:T^{*}\mathbb{T}^{n}\rightarrow\mathbb{R}^{n}$:
@@ 2.4. Linearization of non-degenerate singularities
A singular point $p\in M$ of rank zero of an integrable system $(M,\omega,F=(f_{1},...,f_{n}))$ is non-degenerate if the Hessians $d^{2}f_{1}(p),...,d^{2}f_{n}(p)$ span a so called Cartan subalgebra of the Lie algebra of quadratic forms on $T_{p}M$. We refer to Bolsinov & Fomenko [bolsinov2004integrable] for the precise definition of non-degenerate points of higher rank.
Non-degenerate singularities are in fact linearizable, more precisely, there exists a local normal form which is based on the works of Rüssmann [russmann1964verhalten], Vey [vey1978certains], Colin de Verdiére &\&& Vey [de1979lemme], Eliasson [eliasson1990normal, eliassonphdthesis], Dufour and Molino [dufour1991compactification], Miranda [miranda2014integrable, mirandaphdthesis], Miranda &\&& Zung [miranda2004equivariant], Miranda &\&& Vũ Ngǫc [miranda2005singular], Vũ Ngǫc &\&& Wacheux [vu2013smooth], and Chaperon [chaperon2013normalisation]: there are symplectic local coordinates $(x_{1},...,x_{n},\xi_{1},...,\xi_{n})$ near the singular point $m$ in which $m=(0,...,0)$ corresponds to the origin, the symplectic form assumes the expression $\omega=\sum_{i=1}^{n}dx_{i}\wedge d\xi_{i}$ and there exist functions $q_{1},...,q_{n}$ of $(x_{1},...,x_{n},\xi_{1},...,\xi_{n})$ such that the integrable system $F=(f_{1},...,f_{n})$ satisfies $\{f_{i},q_{j}\}=0$ for all indices $1\leq i,j\leq n$, where $q_{i}$ is one of the following functions:
@@
where $\phi=(x_{1},...,x_{n},\xi_{1},...,\xi_{n})^{-1}$ and $g$ is a diffeomorphism from a small neighborhood of $(0,...,0)$ into another such neighborhood such that $g(0,...,0)=(0,...,0)$. In this paper, we will mostly work with singularities without components of hyperbolic type.
@@ 2.5. Toric systems
The periodicity condition implies that the singularities of toric integrable systems cannot have focus-focus or hyperbolic components, that is, if $m=(0,...,0)$ and $\omega=\sum_{i=1}^{n}dx_{i}\wedge dy_{i}$, then the integrable system is locally in a neighborhood of $m$ of the form
@@
for some $0\leq k\leq n$. Toric integrable systems have connected fibers, see Atiyah [atiyah1982convexity]. In particular, all fibers of a toric system $F$ are diffeomorphic to tori of varying dimensions $\mathbb{T}^{k}$ with $k\in\{0,...,n\}$. Note that this is not the case for general integrable systems.
One of the fundamental theorems of equivariant symplectic geometry, due to Atiyah [atiyah1982convexity] and Guillemin & Sternberg [guillemin1982convexity] says that, if $M$ is compact and connected, the image $F(M)$ is a convex polytope in $\mathbb{R}^{n}$, obtained as the convex hull of the images of the fixed points of the Hamiltonian action of the $n$-torus on $M$. This polytope has the property that if two toric integrable systems are isomorphic then their associated images coincide, possibly up to translations and composition with a matrix in $\text{GL}(n,\mathbb{Z})$. Furthermore, Delzant showed that the convex polytopes obtained as images of toric integrable systems are of a special type: they are simple, rational, and smooth. Such polytopes are called Delzant polytopes and being simple, rational, and smooth means that there are precisely $n$ edges meeting at each vertex, that the slopes of the edges are rational, and that the normal vectors to the facets meeting at each vertex form a basis of the integral lattice. Moreover, Delzant proved that such polytopes are in bijective correspondence with toric integrable systems on compact manifolds in the following sense:

### v2

# Sonja Hohloch  &\&  Pedro Santos

@@ Abstract.
We adapt and generalize McDuff’s method of probes from toric systems to semitoric integrable systems and apply it to study the (non)displaceability properties of the fibers of $3$ examples of semitoric integrable systems.
@@ 1. Introduction
Let $(M,\omega)$ be a closed symplectic manifold. The Arnold conjecture, formulated in the $1960$s, can be stated in many different ways: for instance via the fixed points of a Hamiltonian diffeomorphism $\phi$ or, equivalently, looking at the intersection points of the graph of $\phi$ and the diagonal in $M\times M$. When one endows $M\times M$ with the symplectic form $\omega\oplus-\omega$ the graph of $\phi$ and the diagonal become Lagrangian submanifolds. Therefore the question about the existence of fixed points is turned into a so called Lagrangian intersection problem. This was the approach taken by Floer [Flo88, Flo88a, Flo89] towards the end of the 1980s to prove the Arnold conjecture under certain assumptions. Roughly speaking, for a Lagrangian submanifold $L\subset M$ with $\pi_{2}(M,L)=\{0\}$, Floer developed a homology theory generated by the intersection points of $L\cap\phi(L)$ where $\phi$ is a Hamiltonian diffeomorphism.
@@
Therefore, the interest in (non)displaceability is at least twofold: on the one hand motivated by the computation of Lagrangian Floer homology and on the other hand by symplectic rigidity. Over the years, research on (non)displaceability developed into an area of its own, for instance, just to mention a few, it is essential for Fukaya categories (see Fukaya & Oh & Ohta & Ono [F0009a, F0009b]), it motivated the development of symplectic quasi-states by Entov & Polterovich [EP06], and the development of explicit methods, see for example McDuff [McD11] and Abreu & Borman & McDuff [ABM14], that give sufficient conditions for a Lagrangian torus fiber to be displaceable. This last aspect is in fact the starting point for the present paper, as we will explain now.
Briefly summarized, McDuff’s [McD11] explicit method (called the method of probes) makes use of the Delzant polytope (= momentum polytope) of a symplectic toric manifold. In the world of integrable systems, toric systems are the simplest examples and are classified by their corresponding Delzant polytope. The next natural, more general class of systems are so called semitoric systems (see Definition 2.17) which were classified by Pelayo & Vũ Ngǫc [PV09, PV11] and Palmer & Pelayo & Tang [PPT24] in terms of $5$ invariants. One of these invariants, the so called polytope invariant, is the generalization of the Delzant polytope of a toric system. In the present paper we will generalize the method of probes from toric systems to semitoric systems and apply it to several examples.
@@ 1.1. Our results
Let $(M,\omega,F=(J,H))$ be a semitoric system (see Definition 2.17) on a $4$-dimensional symplectic manifold. Intuitively semitoric systems are more general than toric systems by requiring only an $\mathbb{S}^{1}$-symmetry instead of a $\mathbb{T}^{2}$-symmetry and allowing for focus-focus singularities in addition to the ones appearing in the toric case. As in the toric situation the fibers are connected. A fiber with $k$ focus-focus points is often seen as a torus with $k$ pinches. Alternatively the fiber $F^{-1}(c)$ can be seen as a closed chain of Lagrangian $k$ spheres joint at the poles. If $k=1$ the sphere is immersed, and if $k>1$ the spheres are embedded. Our first result is about the (non)displaceability of focus-focus fibers with more than one focus-focus point:
@@
Figure 1.2. $A(f_{\vec{\epsilon}}(F(M)))$ is the result of applying a suitable integral affine transformation $A$ to the representative of the polytope invariant $f_{\vec{\epsilon}}(F(M))$. Let $0<h<a<\frac{R}{2}$. The square $0,a]\times[0,a]$ fits inside the rectangle $0,R]\times[0,a]$. The rectangle is inside $A(f_{\vec{\epsilon}}(F(M)))$ and does not intersect the cut associated with the focus-focus value $c_{2}$. The rectangle $0,R]\times[0,a]$ represents the symplectic embedding of the product of disks $D^{2}(R)\times D^{2}(a)$. Under these conditions the focus-focus fiber given by $c_{1}$ is displaceable, since we can displace $D^{2}(a)\times D^{2}(a)$ inside $D^{2}(R)\times D^{2}(a)$, due to the fact that $a<\frac{R}{2}$, and we can "torically smoothen" the corner introduced by $c_{1}$.
@@
First, in Section 4.1 we study the coupled spin-oscillator which is a semitoric system $(M,\omega,F)$ on $M=\mathbb{S}^{2}\times\mathbb{R}^{2}$, for which the main result is:
@@
In Section 4.2 we study the coupled angular momenta and its variations. Let $R_{1},R_{2}\in\mathbb{R}^{>0}$ and $t\in[0,1]$ a parameter. Consider the product manifold $M:=\mathbb{S}^{2}\times\mathbb{S}^{2}$ with symplectic form $\omega:=-(R_{1}\omega_{\mathbb{S}^{2}}\bigoplus R_{2}\omega_{\mathbb{S}^{2}})$ where $\omega_{\mathbb{S}^{2}}$ is the standard symplectic form on $\mathbb{S}^{2}$. The coupled angular momenta system is a family of $4$-dimensional completely integrable systems $(M,\omega,F_{t}:=(L,H_{t}))$, with $R_{1}<R_{2}$. The case $R_{1}>R_{2}$ is called the reverse coupled angular momenta and the case $R_{1}=R_{2}$ is called the Kepler problem. For an appropriate range of the parameter $t$ a focus-focus point always exists.
@@
Note that the symplectic rigidity problem is sensitive to the choice of $R_{1}$ and $R_{2}$. For the Kepler problem, concerning the question of (non)displaceability, we get a different behavior:
@@
Let $R_{1}=R_{2}$ and $t$ be in the parameter range such that a focus-focus point exists. There exists a parameter value $t_{0}$ such that:
  • If $t<t_{0}$ there exists only a single nondisplaceable fiber, which is nondisplaceable by symplectomorphisms. In particular the focus-focus fiber is displaceable.
@@
In Section 4.3 we study the (non)displaceability of the fibers of the toric system induced by the octagon $\Delta$ with vertices
@@
as a Delzant polytope. Then we consider the semitoric perturbation of this system, introduced in De Meulenaere & Hohloch [DH21], and study the (non)displaceability properties of the semitoric fibers of this perturbation. Let $(M,\omega,F_{t})$ be the semitoric perturbation of the toric system. For $t=\frac{1}{2}$ the system has two focus-focus fibers with two focus-focus points and otherwise $4$ focus-focus fibers with a single focus-focus point. In Appendix A we show that a representative of the polytope invariant, for $t\neq\frac{1}{2}$, is given by the octagon $\Delta$.
@@
      The fiber over the point $(\frac{3}{2},\frac{3}{2})$ in any representative of the polytope invariant and the focus-focus fibers are nondisplaceable.
@@
      If the image of the focus-focus values in the representative of the polytope invariant given by the octagon are contained in the region $\{y>2\}\cup\{y<1\}$ then the fibers over the points $\{(1,1),(2,1),(\frac{3}{2},\frac{3}{2}),(1,2),(2,2)\}$ are nondisplaceable. All other fibers are displaceable. In particular the focus-focus fibers are displaceable.
@@ 1.2. Structure of the paper
  • In Section 4 we study the (non)displaceability properties of $3$ semitoric systems, namely the coupled spin-oscillator, the coupled angular momenta and the semitoric system induced by the octagon.
@@ 2.2. Integrable systems
Let $(M,\omega)$ be a $2n$-dimensional symplectic manifold, and $\{\cdot\ ,\cdot\}$ be the Poisson bracket induced by $\omega$. An integrable system is a triple $(M,\omega,F)$ where $F:M\rightarrow\mathbb{R}^{n}$ is given by a collection of $n$ real valued smooth functions $f_{1},...,f_{n}$ on $M$ such that:
@@
Two integrable systems $(M,\omega,F)$ and $(M^{\prime},\omega^{\prime},F^{\prime})$ are isomorphic if there exists a diffeomorphism $\psi:M\rightarrow M^{\prime}$ such that $\psi^{*}\omega^{\prime}=\omega$ and a diffeomorphism $g:F(M)\rightarrow F^{\prime}(M^{\prime})$ such that $F^{\prime}\circ\psi=g\circ F$.
@@
A point $m\in M$ in which $X_{f_{1}}(m),...,X_{f_{n}}(m)$ of $T_{m}M$ are linearly dependent is said to be singular or a singularity of $F$. Otherwise we call it regular. The rank of $m$ is given by $\dim\operatorname{span}_{\mathbb{R}}\{X_{f_{1}}(m),...,X_{f_{n}}(m)\}$. If a fiber $F^{-1}(c)$ for some $c\in{\mathbb{R}}^{m}$ contains a singularity we call it singular and otherwise regular. Denote a $k$-dimensional torus briefly by $\mathbb{T}^{k}$. By the Liouville-Arnold-Mineur Theorem, see Arnold [Arn13], a pre-image of a regular value is diffeomorphic to $\mathbb{T}^{n}$. Moreover, $\omega$ vanishes along regular fibers $F^{-1}(c)$, hence $F^{-1}(c)$ is a Lagrangian submanifold, traditionally called a Liouville torus. One often refers to an integrable system $(M,\omega,F)$ as a singular Lagrangian fibration where singular emphasizes that $F$ may have singular fibers of various kinds: for example tori of dimension $m\in\{0,...,n-1\}$, pinched-tori etc.
@@ 2.3. Action-angle coordinates
One important theorem about the symplectic geometry of integrable systems is the existence of so-called action-angle coordinates, see Arnold [Arn63]: each regular fiber, in addition to being diffeomorphic to $\mathbb{T}^{n}$, can be seen as sitting inside of the cotangent bundle $T^{*}\mathbb{T}^{n}$ as the zero section and, in a neighborhood of the fiber, the integrable system can be written in a symplectic normal form $F:T^{*}\mathbb{T}^{n}\rightarrow\mathbb{R}^{n}$:
@@ 2.4. Linearization of non-degenerate singularities
A singular point $p\in M$ of rank zero of an integrable system $(M,\omega,F=(f_{1},...,f_{n}))$ is non-degenerate if the Hessians $d^{2}f_{1}(p),...,d^{2}f_{n}(p)$ span a so called Cartan subalgebra of the Lie algebra of quadratic forms on $T_{p}M$. We refer to Bolsinov & Fomenko [BF04] for the precise definition of non-degenerate points of higher rank.
Non-degenerate singularities are in fact linearizable, more precisely, there exists a local normal form which is based on the works of Rüssmann [Rüs64], Vey [Vey78], Colin de Verdiére &\& Vey [CV79], Eliasson [Eli90, Eli90a], Dufour and Molino [DM91], Miranda [Mir14, Mir03], Miranda &\& Zung [MZ04], Miranda &\& Vũ Ngǫc [MV05], Vũ Ngǫc &\& Wacheux [VW13], and Chaperon [Cha13]: there are symplectic local coordinates $(x_{1},...,x_{n},\xi_{1},...,\xi_{n})$ near the singular point $m$ in which m=(0,,0)m=(0,...,0) corresponds to the origin, the symplectic form assumes the expression $\omega=\sum_{i=1}^{n}dx_{i}\wedge d\xi_{i}$ and there exist functions $q_{1},...,q_{n}$ of $(x_{1},...,x_{n},\xi_{1},...,\xi_{n})$ such that the integrable system $F=(f_{1},...,f_{n})$ satisfies $\{f_{i},q_{j}\}=0$ for all indices $1\leq i,j\leq n$, where $q_{i}$ is one of the following functions:
@@
where $\phi=(x_{1},...,x_{n},\xi_{1},...,\xi_{n})^{-1}$ and $g$ is a diffeomorphism from a small neighborhood of (0,,0)(0,...,0) into another such neighborhood such that g(0,,0)=(0,,0)g(0,...,0)=(0,...,0). In this paper, we will mostly work with singularities without components of hyperbolic type.
@@ 2.5. Toric systems
The periodicity condition implies that the singularities of toric integrable systems cannot have focus-focus or hyperbolic components, that is, if m=(0,,0)m=(0,...,0) and $\omega=\sum_{i=1}^{n}dx_{i}\wedge dy_{i}$, then the integrable system is locally in a neighborhood of $m$ of the form
@@
for some $0\leq k\leq n$. Toric integrable systems have connected fibers, see Atiyah [Ati82]. In particular, all fibers of a toric system $F$ are diffeomorphic to tori of varying dimensions $\mathbb{T}^{k}$ with $k\in\{0,...,n\}$. Note that this is not the case for general integrable systems.
One of the fundamental theorems of equivariant symplectic geometry, due to Atiyah [Ati82] and Guillemin & Sternberg [GS82] says that, if $M$ is compact and connected, the image $F(M)$ is a convex polytope in $\mathbb{R}^{n}$, obtained as the convex hull of the images of the fixed points of the Hamiltonian action of the $n$-torus on $M$. This polytope has the property that if two toric integrable systems are isomorphic then their associated images coincide, possibly up to translations and composition with a matrix in $\text{GL}(n,\mathbb{Z})$. Furthermore, Delzant showed that the convex polytopes obtained as images of toric integrable systems are of a special type: they are simple, rational, and smooth. Such polytopes are called Delzant polytopes and being simple, rational, and smooth means that there are precisely $n$ edges meeting at each vertex, that the slopes of the edges are rational, and that the normal vectors to the facets meeting at each vertex form a basis of the integral lattice. Moreover, Delzant proved that such polytopes are in bijective correspondence with toric integrable systems on compact manifolds in the following sense: